# Stabilizing black-box model selection with the inflated argmax

### v2

@@ 5 Experiments
This definition of stability measures, for trial $j$, the proportion of LOO selected models $\hat{M}_{j}^{\setminus i}$ for $i\in[n]$ that have no overlap with the set of selected models $\hat{M}_{j}$ returned by training on the entire dataset. We also compute the empirical cumulative distribution function (CDF)
@@
where $m^{*}$ is the “true” data generating model in the synthetic experiments. This measure down-weights the accuracy by the number of models returned $|\hat{M}_{j}|$. This metric is used to compare methods in the center column of Figure 3.
@@ 5.1 Model discovery of Lotka-Volterra dynamics
5.1.1 Base algorithm $\mathcal{A}$: Sparse Identification of Nonlinear Dynamical Systems (SINDy)
@@
where $\Theta(u)\in\mathbb{R}^{n\times p}$ ($n$ samples of $p$ functionals) is the library of functionals of $u(t)$ (e.g., polynomial transformations), and $\Xi\in\mathbb{R}^{p\times d}$ are the fitted coefficients. $\Xi$ is assumed to be sparse, so our goal is to find which elements of $\Xi$ should be nonzero, which will immediately give the estimated structure of the governing equations. SINDy finds a sparse solution to $\Xi$ by solving a sparse linear regression problem; see Appendix B for more details on sparse regression.
@@ 5.1.2 Data generation
To construct our library of functions, we selected degree-two polynomials as our covariates, and for one dataset, we visualize the corresponding covariate correlations in Figure 2. As seen in the correlation matrix, the covariate structure contains highly correlated variables (e.g., $u^{(1)}$ and $(u^{(1)})^{2}$), which impacts the model selection stability of SINDy.
In our experiments, we solve the linear regression problem with the STRidge penalization (Rudy et al.,, 2017) as our base algorithm $\mathcal{A}$, as in Fasel et al., (2022).
@@ 5.1.3 Results
We see that across various choices of $\varepsilon$, the inflated argmax provides enhanced stability while at the same time yielding small and accurate sets of selected model(s). This boost in utility-weighted accuracy is due to the adaptivity of the inflated argmax: this method will output a single model as often as possible, whereas for example, top-$k$ will always output $k$ model selections, leading to a comparatively lower utility in many cases. This point is further confirmed in the plot of the median number of models returned vs. worst-case instability in Figure 3.
For a fixed level of worst-case instability empirically seen, we can visualize the empirical CDF to assess the distribution of instability values $\delta_{j}$ across trials $j\in[N]$. We see in Figure 3 that the inflated argmax provides a distribution of instability values much closer to 0 (i.e., higher stability) compared to all other conventional approaches investigated.
In summary, Figure 3 shows that (a) selection methods that return a single model (argmax and ipτ, even with subagging the base algorithm) are less stable than methods that return one or more selected models (i.e., the inflated argmax and top-$k$), (b) the inflated argmax provides higher utility-weighed accuracy scores for similar worst-case stability levels as top-$k$, meaning that the inflated argmax is returning smaller sets containig the correct model more often and (c) the inflated argmax provides smaller sets of returned models compared to top-$k$ for a wide range of worst-case stability values.
@@ 5.2 Graph selection on flow cytometry data
$\hat{\Theta}=\mathop{\!\arg\!\max}_{\Theta}\ \log\det\Theta-\text{tr}(\tilde{\Sigma}\Theta)-\lambda||\Theta||_{1},$ (11)
where $||\cdot||_{1}$ is the entry-wise $l_{1}$ norm, tr()($\cdot$)( ⋅ ) is the trace, $\tilde{\Sigma}$ is the sample covariance matrix, and $\lambda$ is the penalization hyperparameter. In our experiments, we select $\lambda$ via cross-validation, as described in §D.2.1, and hold this quantity fixed across our experiments.
@@
SELECTION LOO AVG. LOO
METHOD INSTABILITY SET SIZE
@@
$\ \mathop{\!\arg\!\max}^{0.02}$ $\circ\ \tilde{\mathcal{A}}_{K,B}$ 0.008 1.58
@@ 6 Discussion
The combination of bagging and the inflated argmax described in this paper can be applied to any black-box model selection procedure and offer stability guarantees described in Theorem 1. We define a notion of stability that is distinct from other works, one that computes the proportion of LOO selected models that are disjoint from the selection made by training on the full data. This is in contrast to, for example, covariate-level stability in linear models.
@@ Appendix B Variable selection in linear models with LASSO
where $\lambda$ is a hyperparameter controlling the sparsity of $\beta$, and the regularization term $R(\beta)=||\beta||_{1}$. Then select variables via
@@ C.1 Data generation
We solve each optimization problem (i.e., fits on the full datasets $\mathcal{D}$, LOO datasets $\mathcal{D}^{\setminus i}$ for $i\in[n]$, and bags $\mathcal{D}^{b}$ for $b\in[B]$ across trials $j\in[N]$) for $\hat{\beta}$ using the optimization function in (17) for $R(\beta)=||\beta||_{1}$, which is equivalent to the LASSO.
@@ C.2.1 Base algorithm $\mathcal{A}$: LASSO
Our learning algorithm $\mathcal{A}$ is the LASSO, which corresponds to solving (17) with $R(\beta)=||\beta||_{1}$, and $\tilde{\mathcal{A}}_{K,B}$ corresponds to a subbagged LASSO with $B=10,000$ bags and $K=25$ training samples per bag.
@@ C.3 Larger sample size ($n=300$)
Figure 9: Models with the top 10 largest weights assigned via subbagging with $B=10,000$ bags for one trial. The y-axis labels correspond to the model weights estimated with 10,000 bags. Covariates within each model (row) that were selected to have a nonzero estimated coefficient by LASSO are shown in black, and covariates that were not selected are shown in white. Only the first 5 covariates are shown since the top 10 models did not select covariates any of the covariates $x^{(6)},x^{(7)},\dots,x^{(200)}$. (Left) Top-10 models using the full dataset where $n=300$. (Right) Top-10 models using a LOO dataset where $n-1=299$.
@@ D.1.1 Data generation
We generate data from the Lotka-Volterra differential equations described in (8) for $(\alpha,\beta,\gamma,\zeta)=(2/3,4/3,1,1)$ for 21 evenly spread time points $t$ in the range $t\in[0,16]$ and initial condition $[1,1]^{\top}$. Solutions of this ordinary differential equation were obtained with the LSODA (Petzold,, 1983) solver. For each $t\in(0,16]$, non-inclusive of the initial condition, we contaminate each $u^{(1)}(t)$ and $u^{(2)}(t)$ with $\mathcal{N}(0,0.05^{2})$ Gaussian noise, where 0.05 is the standard deviation of the measurement noise level.
To obtain sparse solutions of the regression problem and therefore estimate the governing equations, we solve the optimization problem with STRidge (Rudy et al.,, 2017), which iteratively (a) solves the equation in (17) with $R(\beta)=||\beta||^{2}$, corresponding to solving Ridge regression (Hoerl and Kennard,, 1970), and (b) thresholds $\hat{\beta}$ based on $\omega$, where $\omega$ is the minimum magnitude needed for a coefficient in the weight vector. Any coefficient with an estimated magnitude below $\omega$ is set to 0. This process is repeated until a convergence criterion is met.
@@ D.1.2 Cross-validating optimization hyperparameters
D.2 Graph subset selection
D.2.1 Cross-validation to select $\lambda$
@@
Appendix E Stability with varying numbers of bags

### v3

@@ 5 Experiments
This empirical measure of stability measures, for trial $j$, the proportion of LOO selected models $\hat{M}_{j}^{\setminus i}$ for $i\in[n]$ that have no overlap with the set of selected models $\hat{M}_{j}$ returned by training with access to the entire dataset. We also compute the empirical cumulative distribution function (CDF)
@@
[…]
@@ 5.1 Model discovery of Lotka-Volterra dynamics
5.1.1 Base algorithm $\mathcal{A}$: Sparse Identification of Nonlinear Dynamical Systems (SINDy) with Sequentially Thresholded Ridge Regression (STRidge)
@@
where $\Theta(u)\in\mathbb{R}^{n\times p}$ ($n$ samples of $p$ functionals) is the library of functionals of $u(t)$ (e.g., polynomial transformations), and $\Xi\in\mathbb{R}^{p\times d}$ are the fitted coefficients. $\Xi$ is assumed to be sparse, so our goal is to find which elements of $\Xi$ should be nonzero, which will immediately give the estimated structure of the governing equations. SINDy finds a sparse solution to $\Xi$ by solving a sparse linear regression problem; see §B for more details on sparse regression.
@@ 5.1.2 Data generation
To construct our library of functions, we selected degree-two polynomials as our covariates, and for one dataset, we visualize the corresponding covariate correlations in Figure 2. As seen in the correlation matrix, the covariate structure contains highly correlated variables (e.g., $u^{(1)}$ and $(u^{(1)})^{2}$), which impacts the model selection stability of SINDy. In our experiments, we solve the SINDy linear regression problem with STRidge penalization (Rudy et al.,, 2017) as our base algorithm $\mathcal{A}$, as in Fasel et al., (2022).
@@ 5.1.3 Results
We see that across various choices of $\varepsilon$, the inflated argmax provides enhanced stability while at the same time yielding small and accurate sets of selected model(s). This boost in utility-weighted accuracy is due to the adaptivity of the inflated argmax: this method will output a single model as often as possible, whereas for example, top-$k$ will always output $k$ model selections, leading to a comparatively lower utility in many cases. This point is further confirmed in the plot of the median number of models returned vs. worst-case instability in Figure 3. For a fixed level of worst-case instability empirically seen, we can visualize the empirical CDF to assess the distribution of instability values $\delta_{j}$ across trials $j\in[N]$. We see in Figure 3 that the inflated argmax provides a distribution of instability values much closer to 0 (i.e., higher stability) compared to all other conventional approaches investigated.
In summary, Figure 3 shows that (a) selection methods that return a single model (argmax and ipτ, even with subagging the base algorithm) are less stable than methods that return one or more selected models (i.e., the inflated argmax and top-$k$), (b) the inflated argmax provides higher utility-weighed accuracy scores for similar worst-case stability levels as top-$k$, meaning that the inflated argmax is returning smaller sets containig the correct model more often and (c) the inflated argmax provides smaller sets of returned models compared to top-$k$ for a wide range of worst-case stability values. We additionally plot the stability across the parameter $B$ in §E.
@@ 5.2 Graph selection on flow cytometry data
[…]
where $||\cdot||_{1}$ is the entry-wise $l_{1}$ norm, tr()($\cdot$) is the trace, $\tilde{\Sigma}$ is the sample covariance matrix, and $\lambda$ is the penalization hyperparameter. In our experiments, we select $\lambda$ via cross-validation, as described in §F, and hold this quantity fixed across our experiments.
@@
SELECTION METHOD LOO INSTABILITY AVG. LOO SET SIZE
@@
(Ours) $\mathop{\!\arg\!\max}^{0.02}$ $\circ\ \tilde{\mathcal{A}}_{K,B}$ 0.008 1.58
@@ 5.2.1 Results
Though there is no agreed-upon “ground truth” graph connection structure for this real world example, we can compare the identified connections between proteins with those established in the literature. For example, Sachs et al., (2005) visualizes connections between proteins based on scientific evidence, which include PIP3-plcg, praf-pmek, PKA-praf, PIP2-PIP3, PKA-p38, and others, where some of these listed connections are mediated via intermediate proteins. These validated connections are seen in Figure 5. We note, however, that hyperparameter $\lambda$ in (11) largely impacts the accuracy by controlling the number of connections retained in the graph. In our experiment, this parameter was chosen in a purely data-driven way in §F, but a different value may have been chosen with domain expertise.
5.3 Selecting the number of clusters for $\kappa$-means
In this section, we present results on simulated data of an unsupervised $\kappa$-means [^3] In this paper, we use $K$ to represent the number of samples in a bag, $k$ to correspond to the number of models returned by a top-$k$ selection procedure, and $\kappa$ to represent the number of clusters in a clustering problem. clustering (MacQueen,, 1967) example where the task is to identify the number of clusters $\kappa$ in the data. This examples illustrates our algorithm’s use on non-variable selection types of problems. Another example on decision trees can be found in §H. We generate $N=100$ independent datasets, each with sample size $n=30$ with a true number of clusters $\kappa=3$. Further details of our synthetic data generation can be found in §G.
5.3.1 Results
[…]
The leftmost panel shows the empirical distribution of stabilities across methods, with fixed values of $\varepsilon=0.3$ and $k=2$ for argmaxε and top-$k$. Choosing $k=2$ provides 100% stability, meaning there are seemingly only two reasonable choices for the number of clusters $\kappa$. The argmax0.3 showed an empirical worst-case instability of $0.067$ and an average returned set size of 1.18 across LOO trials, which provides high stability with minimal returned models. The middle plot of Figure 6 shows that our approach provides the best utility-weighted accuracy and stability compared to competing methods. In this setting, accuracy is measured based on how often the true $\kappa=3$ is returned in $\hat{M}_{j}$ across all $j=1,\dots,N$ trials). This result shows that the inflated argmax returns sets that are both accurate and as small as possible at each fixed $\varepsilon$ value. This point is further confirmed by looking at the rightmost plot of the median number of returned models versus worst-case instability. Across a wide range of choices of worst-case instabilities, the inflated argmax consistently returns interpretable sets.
Figure 6: Unsupervised $\kappa$-means clustering results. (Left column) Empirical CDF for the stability measures $\delta_{j}$, computed in (5), across $j\in[N]$ for each $\mathcal{S}$. We chose the parameters $k=2,$ and $\varepsilon=0.3$. (Center column) Utility-weighted accuracy, computed in (7), versus the worst-case instability across $\delta_{j}$. (Right column) Median number of models returned across $j$ versus the maximum $\delta_{j}$ across $j\in[N]$ for each $\mathcal{S}$. (Center and right columns) We plot the range of values $k\in[1,2,3]$, and $\epsilon\in[0.005,0.01,0.05,0.1,0.2,0.3,0.5,0.7,0.95,1.0]$.
@@ 6 Discussion
Our framework provides multiple opportunities for further investigation, including exploring the interplay of stable model selection and identifiability, as well as stabilizing model selection without bagging, therefore reducing computation. Moreover, this work primarily focuses on providing theoretical stability guarantees for model selection, but in practice, a practitioner would also be interested in maximizing model performance. Additional future work could explore maximizing performance subject to stability guarantees.
We believe another interesting direction is developing similar theoretical stability results for other measures of stability aside from Definition 2.1. Our definition of stability is relatively weak, reflecting a bare minimum criterion necessary (though not sufficient) for many other notions of stability. Moreover, a limitation of our stability measure is that it may favor larger sets since two returned sets need to overlap by only one model to achieve a score of 1 for that LOO trial. For example, if $|\hat{M}_{j}|$ and $|\hat{M}_{j}^{\setminus i}|$ are large but $|\hat{M}_{j}\ \cap\ M_{j}^{\setminus i}|=1$, the LOO stability for trial $i$ would be 1, which may not reflect what one would colloquially deem as “stable.” An interesting alternative stability measure to theoretically analyze would be one that computes the percent overlap between elements of $\hat{M}_{j}$ and $\hat{M}_{j}^{\setminus i}$.
A natural follow up question to our approach is what a practitioner should do if multiple models are returned. As noted in §5.2.1, top models that substantially overlap can prompt directed follow-up experiments. Moreover, the size of the set of returned models may be a useful indicator of model uncertainty for a practitioner. The decision for how to handle multiple models should be carefully considered and align with the practitioner’s ultimate goals (e.g., model interpretation or predictive accuracy).
The combination of bagging and the inflated argmax described in this paper can be applied to any black-box model selection procedure and offer stability guarantees described in Theorem 1. We define a notion of stability that is distinct from other works, one that computes the proportion of LOO selected models that are disjoint from the selection made by training with access to the full data. This is in contrast to, for example, covariate-level stability in linear models.
@@ Appendix B Variable selection in linear models with LASSO
[…]
@@ C.1 Data generation
[…]
@@ C.2.1 Base algorithm $\mathcal{A}$: LASSO
[…]
@@ C.3 Larger sample size ($n=300$)
Figure 10: Models with the top 10 largest weights assigned via subbagging with $B=10,000$ bags for one trial. The y-axis labels correspond to the model weights estimated with 10,000 bags. Covariates within each model (row) that were selected to have a nonzero estimated coefficient by LASSO are shown in black, and covariates that were not selected are shown in white. Only the first 5 covariates are shown since the top 10 models did not select covariates any of the covariates $x^{(6)},x^{(7)},\dots,x^{(200)}$. (Left) Top-10 models using the full dataset where $n=300$. (Right) Top-10 models using a LOO dataset where n1=299n-1=299.
@@ D.1.1 Data generation
We generate data from the Lotka-Volterra differential equations described in (8) for $(\alpha,\beta,\gamma,\zeta)=(2/3,4/3,1,1)$ for 21 evenly spread time points $t$ in the range $t\in[0,16]$ and initial condition $1,1]^{\top}$. Solutions of this ordinary differential equation were obtained with the LSODA (Petzold,, 1983) solver. For each $t\in(0,16]$, non-inclusive of the initial condition, we contaminate each $u^{(1)}(t)$ and $u^{(2)}(t)$ with $\mathcal{N}(0,0.05^{2})$ Gaussian noise, where 0.05 is the standard deviation of the measurement noise level.
[…]
@@ D.1.2 Cross-validating optimization hyperparameters
Appendix E Stability with varying numbers of bags: Lotka-Volterra example
Appendix F Graph subset selection
F.1 Cross-validation to select $\lambda$
@@
Appendix G $\kappa$-means clustering
G.1 Data generation
We randomly generate $N=100$ independent datasets in $\mathbb{R}^{n\times 2}$ with $n=30$ total data points per dataset and $\kappa=3$ distinct clusters. To generate the three clusters, we generate $n_{1}=5$ samples from $\mathcal{N}\left([1.5,1.5]^{\top},I_{2}\right)$, $n_{2}=5$ samples from $\mathcal{N}([3.5,3.5]^{\top},I_{2})$, and $n_{3}=20$ samples from $\mathcal{N}([2.5,2.5]^{\top},0.3^{2}I_{2})$, where $\mathcal{N}$ denotes a Gaussian distribution, and $I_{2}$ is a $2\times 2$ identity matrix. An example of one of our generated datasets is shown in Figure 15.
Figure 15: Example clustering dataset where the sample size is $n=30$ and the number of clusters is $\kappa=3$.
G.2 Base algorithm $\mathcal{A}$: the “elbow” plot from $\kappa$-means
A common technique in unsupervised clustering for selecting the number of clusters $\kappa$ is via the “elbow” of the plot of the sum of squared distances (SSD) vs. $\kappa$. We deterministically select $\kappa$ as described in Algorithm 2, which selects the number of clusters by some hyperparameter for the $\kappa$ value that leads to the minimum observable slope of the SSD vs. $\kappa$ plot. In the SSDκ calculation in Algorithm 2, $S_{i}$ is the set of points in cluster $i$, and $\mu_{i}$ is the cluster center for cluster $i$ selected via $\kappa$-means. In our generated example, the maximum number of clusters $M=29$, and the slope tolerance $\omega=5$. Additionally, for each run of $\kappa$-means, the clusters are randomly initialized.
Algorithm 2 Computing the number of clusters
0: Maximum number of clusters $M$, dataset $X$, slope tolerance $\omega$
for $\kappa=1,\ldots,M$ do
  Compute the sum of squared distances (SSD):
[…]
  Compute the slope of SSD curve using the current SSDκ and the previous SSDκ-1:
$\textrm{slope}_{\kappa}=\frac{\textrm{SSD}_{\kappa}-\textrm{SSD}_{\kappa-1}}{\kappa-(\kappa-1)}$
[…]
   output Number of clusters $\kappa$
  end if
end for
 Number of clusters $M$
Appendix H Decision tree classification on single cell transcriptomics data
We investigate a model selection problem involving deciding the structure of a decision tree for multiclass classification. In this setting, we only measure two decision trees as distinct from each other if the trees split on the same features for a given split. Therefore, we do not consider the exact split value for continuous features in deciding whether two selected models are the same.
We explore the decision tree stability with data from single cell transcriptomics to classify stem cell fate based on gene expression. Mouse embryonic stem cells were sequenced for their gene expression and subsequently labeled as three different cell types: embryonic stem cells (ESC), epiblast-like intermediate cells (EPI), and neural progenitor cells (NPC) (Veleslavov and Stumpf,, 2020). The cell fate label is noted to be assigned by clustering, so there is uncertainty in the true cell type label.
The dataset contains 547 cells and 96 gene expression values per cell. To explore stability in a setting where the sample size is much lower than the feature dimensionality, which is a common problem in genetics, we randomly subsample the number of cells to 50 and compute decision tree stabilities. As shown in Figure 16, gene expression is high correlation among genes, which may lead to instabilities in the fitted decision tree for small perturbations in the training data.
Figure 16: Empirical correlation matrix of gene expression across 50 cells from data in Veleslavov and Stumpf, (2020). Rows and columns that show a null value correspond to genes that showed no variability in expression across the 50 cells.
The base algorithm $\mathcal{A}$ corresponds to a multiclass classification tree. Broadly, decision trees create a sequence of splits on features that sort the training data samples into nodes, where a particular split is chosen to minimize the impurity of the samples in the two split nodes. Greedy algorithms are typically implemented to sequentially choose the locally best split at each node. More information on decision trees can be found in Breiman et al., (1984). Additionally, Xin et al., (2022) formulates an algorithm for finding the full Rashomon set for decision trees, which can be used as an alternative base algorithm to rank decision tree structures.
For this classification problem, we choose to measure dissimilarities among samples in a node via Gini impurity, which can be written as
$\sum_{d=1}^{D}p_{ld}(1-p_{ld})$ (26)
for $d\in[D]$ classes at node $l$, where the proportion of class $d$ in node $l$ is
$p_{ld}=\frac{1}{n_{l}}\sum_{y\in Q_{l}}\mathbb{I}\{y=d\},$ (27)
and $Q_{l}$ denotes the split of node $l$ based on feature index $f$ and split value $t_{l}$:
$Q^{\text{left}}_{l}$ $=\{(x,y)|\ x_{f}\leq t_{l}\},$ (28)
$Q^{\text{right}}_{l}$ $=Q_{l}\setminus Q_{l}^{\text{left}}.$ (29)
Our experiments use the default hyper-parameters in the sklearn.tree.DecisionTreeClassifier function.
H.1 Results
Since we have one dataset of gene expression across cells, $N=1$ in this example. Table 4 shows the stability results and average number of models returned across LOO trials. Notably, ipτ is absent from this table because this selection method is specific to variable selection, which is a special class of model selection.
Based on Table 4, significant instabilities are seen for both the $\mathop{\!\arg\!\max}\circ\ \mathcal{A}$ and $\mathop{\!\arg\!\max}\circ\ \tilde{\mathcal{A}}_{K,B}$ model selection methods, while top-2 $\circ\ \tilde{\mathcal{A}}_{K,B}$ provides enhanced stability. For the inflated argmax, we chose $\varepsilon=0.03$ so that the stability matched that of the stability of top-2 in order for us to directly compare the number of returned models. For a fixed level of instability $\delta=0.12$, the inflated argmax returns a smaller number of models on average across LOO trials, showing that our method adaptively returns the smallest set of models based on the underlying uncertainty in the selection.
For $\varepsilon=0.04$, we obtain $0.04$ instability, corresponding to a high degree of stability with a LOO average of 2.88 returned models. We visualize the top 3 selected models across bootstrap samples in Figure 17. From these trees, we can see that the first and second most common selected trees across bootstrap samples are completely disjoint, potentially indicating strong correlations among the features for both trees. However, the third most commonly selected tree differs from the top tree by only one split, suggesting there may be a high correlation between the gene expressions of Lin28a and Nanog. Future investigations can focus on disentangling the associations among the genes selected in the top tree versus the second most selected tree to elucidate their associations with the cell fates.
SELECTION METHOD LOO INSTABILITY AVG. LOO SET SIZE
$\mathop{\!\arg\!\max}\circ\ \mathcal{A}\ \ \ \ \$0.40 1.00
$\mathop{\!\arg\!\max}\circ\ \tilde{\mathcal{A}}_{K,B}$ 0.24 1.00
top-2 $\circ\ \tilde{\mathcal{A}}_{K,B}$ 0.12 2.02
(Ours) $\mathop{\!\arg\!\max}^{0.03}$ $\circ\ \tilde{\mathcal{A}}_{K,B}$ 0.12 1.74
Table 4: Table of stability and the average number of models returned across LOO trials for selecting the structure of the multi-class decision tree for a gene expression dataset from Veleslavov and Stumpf, (2020). In this experiment, $\mathcal{A}$ is a classification decision tree, and $\tilde{A}_{K,B}$ is a bootstrapped classification decision tree with $B=10,000$ and $K=45$. In computing the returned models, if there is an exact tie in the top models, all ties are returned, which explains why top-2 returns slightly more than 2 models on average across the LOO fits.
Figure 17: Visualization of the decision tree structures for the top 2 trees chosen by bootstrap sampling the data for $K=45$ samples per bag across $B=10,000$ bags. Differences between the top and second tree are indicated in blue in the second tree, and the differences between the first and third trees are indicated in red. The names of the genes that determine the split at each node are shown in ovals, and the class labels NPC, ESC, and EPI are listed for each leaf. The top 1 decision tree was selected by 4.74% of bags, the top 2 decision tree was selected by 2.80% of bags, and the third decision tree was selected by 2.24% of bags. The exact values used for the splits at each node were not considered in our stability results.